OpenStudy (anonymous):

Solve 9c + 4 = -23. (1 point)

OpenStudy (anonymous):

-3 3 -2.1 4

OpenStudy (anonymous):

a is the right answer or -3

OpenStudy (texaschic101):

9c + 4 = -23 --- subtract 4 from both sides 9c = -23 - 4 9c = - 27 -- divide both sides by 9 c = -27/9 c = -3 Just in case you wanted to see how to get the answer

OpenStudy (anonymous):

Solve 3x - x + 2 = 12. (1 point) -5 4 7 5 @texaschic101

OpenStudy (anonymous):

could you help with this one to

OpenStudy (texaschic101):

3x - x + 2 = 12 2x + 2 = 12 --- subtract 2 from both sides 2x = 12 - 2 2x = 10 -- divide both sides by 2 x = 5

OpenStudy (anonymous):

thank you last one . Solve y/ 4 + 5 = 15. (1 point) 80 -80 40 -40 @texaschic101

OpenStudy (anonymous):

Solve 2(x + 1) + 4 = 8. (1 point) 1 6.5 4 3

OpenStudy (texaschic101):

y/4 + 5 = 15 --- subtract 5 from both sides y/4 = 15 - 5 y/4 = 10 --- multiply both sides by 4 y = 40 ============== 2(x + 1) + 4 = 8 -- distribute through the parenthesis 2x + 2 + 4 = 8 2x + 6 = 8 -- subtract 6 from both sides 2x = 8 - 6 2x = 2 -- divide both sides by 2 x = 1

OpenStudy (anonymous):

thank you @texaschic101

OpenStudy (texaschic101):

sure thing :)